# Enhancing Adversarial Transferability with Adversarial Weight Tuning

### v2

@@ Introduction
From a practical perspective, one particularly concerning aspect of AEs is their ability to transfer between different models (Ma et al. 2023; Zhang et al. 2023), allowing attackers to craft effective perturbations without direct access to the target model’s architecture or parameters (Dong et al. 2018; Wang et al. 2021b; Wang and He 2021). Recently, researchers have proposed a variety of transferable attack methods aiming to maximize the success rate of these attacks and avoid over-fitting to the surrogate models (Dong et al. 2018, 2019). These methods leverage different strategies, such as input transformation by applying data augmentation (Xie et al. 2019), model ensemble that exploits the outputs of multiple models in loss or logits (Liu et al. 2016) and gradient rectification with momentum (Dong et al. 2018; Lin et al. 2019), to generate transferable AEs across diverse architectures of the surrogate model. Although these methods have presented better performance, there are no clear explanations for the factors behind transferability.
To explain what factors have resulted in the different transferability of different models, Wu et al. (Wu et al. 2018) explored how adversarial transferability relies on the model-specific factors empirically and found that local non-smoothness of the loss surface harms the transferability of generated AEs. Next, Yang et al. (Yang et al. 2021) first theoretically linked two key factors, i.e., model smoothness and gradient similarity, with the lower bound of AEs transferability. With this intuition, researchers have proposed to finetune the attacker’s surrogate model to enhance its smoothness (Zhu et al. 2022; Wu, Luo, and Wunsch II 2024), since AEs generated on smoothed surrogate models have shown better transferability. Nevertheless, these methods all assume that attackers can obtain extra data with a similar distribution of the target models’ training data, and time-consuming tuning on these large-scale datasets introduces computation overhead as well. Different from these model-based attacks, recent works (Ge et al. 2023; Qiu et al. 2024) can achieve state-of-the-art (SOTA) performance on transferability by optimizing the flat local maxima (minima) of AEs, as illustrated (from black to blue dot) in Figure 1.
@@
To this end, we rethink the property of transferable AEs and develop a novel formulation from the perspective of generalization to measure transferability. Building upon the insights from this formulation, we first analyze the generalization of AEs across models with different architectures and prove that we can always find a local perturbation to mitigate the gap between surrogate and target models. Further, with this insights, we observe inner connections between model smoothness and flat local maxima, which both contribute to the transferability of AEs. Based on this, we then propose a new adversarial attack algorithm, Adversarial Weight Tuning (AWT), which adaptively adjusts the parameters of the surrogate model using generated AEs to optimize the flat local maxima and model smoothness simultaneously, without the need for extra data. As shown in Figure 1, better transferability (orange dot) can be achieved through input and parameter space. We evaluate AWT on a variety of models with traditional experiments and the metric derived from our proposed formulation. Our contributions are as follows:
@@
Gradient-based methods are among the most common approaches to generating AEs and improving their transferability. These methods basically involve crafting perturbations based on the gradients of the loss function with respect to the input. One of the earliest and most influential gradient-based methods is the Fast Gradient Sign Method (FGSM) (Goodfellow, Shlens, and Szegedy 2014), which adds a small perturbation to the input in the direction of the sign of the gradient. Subsequent improvements, such as the Iterative FGSM (Kurakin, Goodfellow, and Bengio 2016) and the Momentum-based FGSM (Dong et al. 2018; Lin et al. 2019), iteratively refine the perturbation to achieve higher success rates and better transferability. Later, researchers proposed methods to improve the transferability of AEs by leveraging additional gradient information. For example, the Virtual Momentum Iterative FGSM (VMI) (Wang and He 2021) and the Ensemble Momentum Iterative FGSM (EMI) (Wang et al. 2021b) use virtual and ensemble gradients, respectively, to enhance the transferability of AEs. More recently, the connection between flat local maxima and generalization (Keskar et al. 2016; Foret et al. 2020) have inspired researchers to generate AEs at a flat local region of the loss function, making them more likely to generalize across models. For instance, Qin et al. (Qin et al. 2022) seek to minimize the loss of AEs at a region with low loss value by adding reverse perturbations before calculating gradients. Ge et al. and Qiu et al. (Ge et al. 2023; Qiu et al. 2024) introduce approximated Hessian-vector products and aggregate the gradients of randomly sampled data around AEs.
@@
In addition to gradient-based and model-based methods, researchers have explored other techniques to improve the transferability of AEs. One approach is to apply input transformations, such as data augmentation, to the surrogate model during training. Techniques like Diverse Input (DIM) (Xie et al. 2019) and Translation-invariant method (TIM) (Dong et al. 2019) have been shown to improve the transferability of AEs by making the AEs more robust to data transformations. Another line of research focuses on optimizing AEs with the aggregated gradients from the ensemble of surrogate models (Liu et al. 2016; Gubri et al. 2022). With augmentation of surrogate models, these methods aim to create AEs that are able to fool a wider range of target models with significantly different architectures.
@@
For practical implementation, we incorporate previous gradient-based techniques to achieve flat local maxima in both input and parameter spaces, thereby improving the generalization of AEs across different models. The complete algorithm for generating AEs is detailed in Algorithm 1, which also ensures a low $\|\nabla_{\theta^{s}}\ell(x^{\prime},y;\theta)\|_{2}$ by optimizing samples around $x^{\prime}$.
@@
Compared Methods. As for baselines, we employ eight popular gradient-based iterative adversarial attacks: MI (Dong et al. 2018), Nesterov Accelerated FGSM (NI) (Lin et al. 2019), VMI (Wang and He 2021), EMI (Wang et al. 2021b), and Robust Adversarial Perturbations (RAP) (Qin et al. 2022), Probabilistic Gradient Noise (PGN) (Ge et al. 2023) and Neighborhood Conditional Sampling (NCS) (Qiu et al. 2024).
@@
Integrated with Other Attacks. Having demonstrated the effectiveness of the AWT (only model tuning) method on both CNN- and Transformer-based models, we now investigate its impact when integrated with other attacks which also follows the approach taken in previous works (Qiu et al. 2024). AWT can further enhance the adversarial transferability of other attacks without any conflict or overlap, effectively complementing each other. Specifically, we integrate AWT with input transformation-based attacks, such as DIM (Xie et al. 2019), TIM (Dong et al. 2019), SIM (Lin et al. 2019), and Admix (Wang et al. 2021a), and evaluate the ASR on a variety of target models. Table 3 presents the ASR of these attacks when integrated with AWT, using RN50 and RN101 as the surrogate models. The best results are highlighted in bold. As shown in Table 3, integrating AWT with existing attack methods consistently improves the ASRs across all target models. For instance, when integrated with DIM, AWT increases the ASR from 43.3% to 48.5% on the Inc-v4 model, and from 55.2% to 58.1% on the Inc-v3 model when using RN50 as the surrogate model. Similarly, when integrated with TIM, AWT improves the ASR from 36.9% to 42.9% on Inc-v3 when using RN101 as the surrogate model. The improvements are substantial across all attack methods. For example, when integrated with Admix, AWT increases the ASR from 62.3% to 66.2% on Inc-v3 when using RN50 as the surrogate model. The integration of AWT with other attacks demonstrates its versatility and ability to enhance the transferability of AEs.
@@
Further, with the evidence above, we use our proposed metric to evaluate nine attacks under different parameter perturbations with RN50 and RN101 and the results are shown in Table 4. We can observe that AWT can achieve the lowest metric value among all attacks, suggesting that it generates AEs with better transferability. This finding aligns with the results reported earlier in Table 1 and Table 2. However, we can also notice there are discrepancies between experimental results and our metric. For instance, PGN and NCS tend to perform better transferability than other attacks but they sometimes have larger metric values than attacks than VMI and VNI. This can be attributed to the insufficient sampling of our metric, leading to significant errors. Therefore, it is crucial to acknowledge that the transferability of AEs in real-world settings can be significantly more complex due to the diversity of model architectures. Given these complexities, our proposed metric, while effective in certain contexts, may not fully capture the intricacies of transferability in practice. (Limitations) (1) The metric assumes a certain level of similarity between the surrogate model and the target model with the same architecture, which might not always hold in practical applications. (2) Even if two models share the same architecture, differences in their training data or hyperparameter settings can lead to variations in the transferability of AEs (Wu et al. 2018), which have not been considered in the proposed metric.

### v3

@@ Introduction
From a practical perspective, one particularly concerning aspect of AEs is their ability to transfer between different models (Ma et al. 2023; Zhang et al. 2023), allowing attackers to craft effective perturbations without direct access to the target model’s architecture or parameters (Dong et al. 2018; Wang et al. 2021c; Wang and He 2021). Recently, researchers have proposed a variety of transferable attack methods aiming to maximize the success rate of these attacks and avoid over-fitting to the surrogate models (Dong et al. 2018, 2019). These methods leverage different strategies, such as input transformation by applying data augmentation (Xie et al. 2019), model ensemble that exploits the outputs of multiple models in loss or logits (Liu et al. 2016) and gradient rectification with momentum (Dong et al. 2018; Lin et al. 2019), to generate transferable AEs across diverse architectures of the surrogate model. Although these methods have presented better performance, there are no clear explanations for the factors behind transferability.
To explain what factors have resulted in the different transferability of different models, Wu et al. (Wu et al. 2018) explored how adversarial transferability relies on the model-specific factors empirically and found that local non-smoothness of the loss surface harms the transferability of generated AEs. Next, Yang et al. (Yang et al. 2021) first theoretically linked two key factors, i.e., model smoothness and gradient similarity, with the lower bound of AEs transferability. With this intuition, researchers have proposed to finetune the attacker’s surrogate model to enhance its smoothness (Zhu et al. 2022; Wu, Luo, and Wunsch II 2024; Zhang et al. 2024), since AEs generated on smoothed surrogate models have shown better transferability. Nevertheless, these methods all assume that attackers can obtain extra data with a similar distribution of the target models’ training data, and time-consuming tuning on these large-scale datasets introduces computation overhead as well. Different from these model-based attacks, recent works (Ge et al. 2023; Qiu et al. 2024) can achieve state-of-the-art (SOTA) performance on transferability by optimizing the flat local maxima (minima) of AEs, as illustrated (from black to blue dot) in Figure 1.
@@
To this end, we rethink the property of transferable AEs and develop a novel formulation from the perspective of generalization to measure transferability. Building upon the insights from this formulation, we first analyze the generalization of AEs across models with different architectures and prove that we can always find a local perturbation to mitigate the gap between surrogate and target models. Further, with these insights, we observe inner connections between model smoothness and flat local maxima, which both contribute to the transferability of AEs. Based on this, we then propose a new adversarial attack algorithm, Adversarial Weight Tuning (AWT), which adaptively adjusts the parameters of the surrogate model using generated AEs to optimize the flat local maxima and model smoothness simultaneously, without the need for extra data. As shown in Figure 1, better transferability (orange dot) can be achieved through input and parameter space. We evaluate AWT on a variety of models with traditional experiments and the metric derived from our proposed formulation. Our contributions are as follows:
@@
Gradient-based methods are among the most common approaches to generating AEs and improving their transferability. These methods basically involve crafting perturbations based on the gradients of the loss function with respect to the input. One of the earliest and most influential gradient-based methods is the Fast Gradient Sign Method (FGSM) (Goodfellow, Shlens, and Szegedy 2014), which adds a small perturbation to the input in the direction of the sign of the gradient. Subsequent improvements, such as the Iterative FGSM (Kurakin, Goodfellow, and Bengio 2016) and the Momentum-based FGSM (Dong et al. 2018; Lin et al. 2019), iteratively refine the perturbation to achieve higher success rates and better transferability. Later, researchers proposed methods to improve the transferability of AEs by leveraging additional gradient information. For example, the Virtual Momentum Iterative FGSM (VMI) (Wang and He 2021) and the Ensemble Momentum Iterative FGSM (EMI) (Wang et al. 2021c) use virtual and ensemble gradients, respectively, to enhance the transferability of AEs. More recently, the connection between flat local maxima and generalization (Keskar et al. 2016; Foret et al. 2020) have inspired researchers to generate AEs at a flat local region of the loss function, making them more likely to generalize across models. For instance, Qin et al. (Qin et al. 2022) seek to minimize the loss of AEs at a region with low loss value by adding reverse perturbations before calculating gradients. Ge et al. and Qiu et al. (Ge et al. 2023; Qiu et al. 2024) introduce approximated Hessian-vector products and aggregate the gradients of randomly sampled data around AEs.
@@
In addition to gradient-based and model-based methods, researchers have explored other techniques to improve the transferability of AEs (Wang et al. 2022, 2021a, 2024a, 2024b). One approach is to apply input transformations, such as data augmentation, to the surrogate model during training. Techniques like Diverse Input (DIM) (Xie et al. 2019) and Translation-invariant method (TIM) (Dong et al. 2019) have been shown to improve the transferability of AEs by making the AEs more robust to data transformations. Another line of research focuses on optimizing AEs with the aggregated gradients from the ensemble of surrogate models (Liu et al. 2016; Gubri et al. 2022). With augmentation of surrogate models, these methods aim to create AEs that are able to fool a wider range of target models with significantly different architectures.
@@
For practical implementation, we incorporate previous gradient-based techniques (Qiu et al. 2024; Ge et al. 2023) (flat local maxima of input space) to achieve flat local maxima in both input and parameter spaces, thereby improving the generalization of AEs across different models. The complete algorithm for generating AEs is detailed in Algorithm 1, which also ensures a low $\|\nabla_{\theta^{s}}\ell(x^{\prime},y;\theta)\|_{2}$ by optimizing samples around $x^{\prime}$.
@@
Compared Methods. As for baselines, we employ eight popular gradient-based iterative adversarial attacks: MI (Dong et al. 2018), Nesterov Accelerated FGSM (NI) (Lin et al. 2019), VMI (Wang and He 2021), EMI (Wang et al. 2021c), and Robust Adversarial Perturbations (RAP) (Qin et al. 2022), Probabilistic Gradient Noise (PGN) (Ge et al. 2023) and Neighborhood Conditional Sampling (NCS) (Qiu et al. 2024).
@@
Integrated with Other Attacks. Having demonstrated the effectiveness of the AWT (only model tuning) method on both CNN- and Transformer-based models, we now investigate its impact when integrated with other attacks which also follows the approach taken in previous works (Qiu et al. 2024). AWT can further enhance the adversarial transferability of other attacks without any conflict or overlap, effectively complementing each other. Specifically, we integrate AWT with input transformation-based attacks, such as DIM (Xie et al. 2019), TIM (Dong et al. 2019), SIM (Lin et al. 2019), and Admix (Wang et al. 2021b), and evaluate the ASR on a variety of target models. Table 3 presents the ASR of these attacks when integrated with AWT, using RN50 and RN101 as the surrogate models. The best results are highlighted in bold. As shown in Table 3, integrating AWT with existing attack methods consistently improves the ASRs across all target models. For instance, when integrated with DIM, AWT increases the ASR from 43.3% to 48.5% on the Inc-v4 model, and from 55.2% to 58.1% on the Inc-v3 model when using RN50 as the surrogate model. Similarly, when integrated with TIM, AWT improves the ASR from 36.9% to 42.9% on Inc-v3 when using RN101 as the surrogate model. The improvements are substantial across all attack methods. For example, when integrated with Admix, AWT increases the ASR from 62.3% to 66.2% on Inc-v3 when using RN50 as the surrogate model. The integration of AWT with other attacks demonstrates its versatility and ability to enhance the transferability of AEs.
@@
Further, with the evidence above, we use our proposed metric to evaluate nine attacks under different parameter perturbations with RN50 and RN101 and the results are shown in Table 4. We can observe that AWT can achieve the lowest metric value among all attacks, suggesting that it generates AEs with better transferability. This finding aligns with the results reported earlier in Table 1 and Table 2. However, we can also notice there are discrepancies between experimental results and our metric. For instance, PGN and NCS tend to perform better transferability than other attacks but they sometimes have larger metric values than attacks than VMI and VNI. This can be attributed to the insufficient sampling of our metric, leading to significant errors. Therefore, it is crucial to acknowledge that the transferability of AEs in real-world settings can be significantly more complex due to the diversity of model architectures. Given these complexities, our proposed metric, while effective in certain contexts, may not fully capture the intricacies of transferability in practice. (Limitations) (1) The metric assumes a certain level of similarity between the surrogate model and the target model with the same architecture, which might not always hold in practical applications. (2) Even if two models share the same architecture, differences in their training data or hyperparameter settings can lead to variations in the transferability of AEs (Wu et al. 2018), which have not been considered in the proposed metric. (3) Note that AWT has introduced extra computational cost with two backward for calculating parameter gradients during each iteration while different samples do need different surrogate parameters to achieve better optimal. In practice, attackers should balance the trade-off between computational cost and effectiveness.
@@ Conclusion
Acknowledgement
This work was partly supported by the National Key Research and Development Program of China under No. 2022YFB3102100, NSFC under No. U244120033, U24A20336, 62172243 and 62402425, and the Zhejiang Provincial Natural Science Foundation under No. LD24F020002.